OpenStudy (anonymous):

A map has a scale of 1 cm : 20 km. Two cities are 2.5 cm apart on the map. To the nearest tenth of a kilometer, what is the actual distance corresponding to the map distance?

OpenStudy (anonymous):

1/20 = 2.5/x cross multiply x= 20 x 2.5

OpenStudy (anonymous):

1cm - >20 Km 2.5 -> X The find x

OpenStudy (anonymous):

x= 50Km

OpenStudy (anonymous):

thanks for the help c: